OpenStudy (help108):

Mona solved an equation incorrectly, as shown below: Step 1: 5x = 30 Step 2: x = 30 − 5 Step 3: x = 25 Which statement best explains why Step 2 is incorrect in Mona's solution? She did not add 5 to 30. She did not divide 30 by 5. She did not multiply 30 by 5. She did not multiply 30 five times.

OpenStudy (photon336):

This is wrong because to get the answer you need to divide both sides by 5

OpenStudy (anonymous):

She did not divide 30 by 5. One has to divide each side of the equation by 5 to get rid of the 5 in front of the x term.

OpenStudy (help108):

so B? @Photon336

OpenStudy (anonymous):

Yes.

OpenStudy (help108):

Thanks

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (help108):

An equation is shown below: 4(2x − 3) = 1 Which of the following correctly shows the first two steps to solve this equation? Step 1: 6x − 7 = 1; Step 2: 6x = 8 Step 1: 6x + 1 = 1; Step 2: 6x = 0 Step 1: 8x − 12 = 1; Step 2: 8x = 13 Step 1: 8x − 3 = 1; Step 2: 8x = 4

OpenStudy (help108):

@robtobey

OpenStudy (anonymous):

Step 1: 8x − 12 = 1; Step 2: 8x = 13 The idea is to move all of the integers to the RHS and then divide each side of the equation by the coefficient of x in order to solve for x.